# Differentially Private Knowledge Distillation via Synthetic Text Generation

### v1

@@ 4.3 Privacy Analysis
Proof. Our goal is for our algorithm to guarantee $(\epsilon,\delta)$-DP for $D$, which is the text content of each data sample. After step 1, our method produces a teacher model $T$ that is $(\epsilon,\delta)$-DP. Steps 2 and 3 make use of the post-processing property of DP. In step 2, sampling from $T$ incurs no extra privacy loss due to the post-processing property. However, one subtle, but crucial, detail is the potential privacy leakage from the control codes $C$. We limit the control codes to only contain categorical information of a data sample, not text contents. However, the categorical distribution in the original dataset may be private information. In our work, we follow the same procedure as Yue et al. (2022) by ignoring this low-cost privacy loss and using the exact categorical distribution.
For step 3, since $\tilde{D}$ and $T$ are $(\epsilon,\delta)$-DP with respect to $D$, utilizing a non-private optimizer to minimize Equation 1 will result in $S$ being $(\epsilon,\delta)$-DP due to the post-processing property. ∎
@@ 5 Experiments
GPT2-Large DP-SGD 1.0 - 26.84 21.91
DistilGPT2 Zero-shot 0 - 65.46 57.42
DistilGPT2 DP-SGD 2.0 - 48.12 41.80
DistilGPT2 DPKD 2.0 GPT2-Large 46.16 41.20
DistilGPT2 DP Syn Data 2.0 GPT2-Large 61.84 49.26
DistilGPT2 Ours 2.0 GPT2-Large 44.15 32.43
@@
5.1 Experimental Setup
We experimentally evaluated the privacy-utility tradeoff of our framework with the following setup. We used DistilGPT2 Sanh et al. (2019) and GPT2-Large Radford et al. (2019) as our student and teacher model, respectively. GPT2-Large contains 774M parameters while DistilGPT2 contains 82M parameters, so the teacher model is about 9.5 times larger than the student. Our implementation of fine-tuning language models with DP and knowledge distillation utilizes the dp-transformers library Inan et al. (2022), which combines the NLP library Huggingface Wolf et al. (2019) with the DP library Opacus Yousefpour et al. (2021).
We experimented with the Yelp Open dataset  [^1] https://www.yelp.com/dataset, which contains review text data on businesses. We follow a similar experimental setup as Yue et al. (2022) by using the review stars and business category as the control codes. These control codes are constructed as "Business Type: Restaurant |normal-||| Review Stars: 3.0" and are prepended to each sample. The top 10 frequent business categories are sampled and reviews with no ratings are removed. The resulting data split is $1.9$M reviews for training, $5000$ for validation, and $5000$ for testing. Each data sample is set to a sequence length of $128$.
We compare our method against three baselines: (1) a student model fine-tuned directly with DP-SGD; (2) a student model fine-tuned using DPKD Mireshghallah et al. (2022); (3) a student model fine-tuned only on DP synthetic data (DP Syn Data), i.e. $\lambda=0$ in Eq. 1, which follows the framework of Yu et al. (2023a). The training hyperparameters used for our method and the baselines with the teacher and student models are shown in table 1. We set the privacy budget to $\epsilon=2$, which is a strong privacy guarantee. We follow the standard practice of setting $\delta=1/N$ where $N$ is the size of the dataset.
For the generation of synthetic text, we use the truncation decoding strategies of top-$k$ Fan et al. (2018) and top-$p$ Holtzman et al. (2019) sampling using $k=50$ and $p=0.9$. The teacher model generates 400K synthetic samples using control codes from the Yelp training dataset and 200K synthetic samples from Big Patent. The maximum sequence length of each generated sequence is $128$. We follow the methodology of Kurakin et al. (2023) for training on DP synthetic data by splitting the Yelp synthetic samples into $396$K for the training set and $4$K for the validation. Likewise, 198K for the training set and 2K for the validation with the Big Patent synthetic examples. For both our method and DPKD, we set $t=1.0$ and $\lambda=0.4$ for Equation 1. One minor technicality of our method is that because our teacher model was trained on control codes, its output distribution on the initial context vectors will be skewed toward the types of control codes. Hence, we ignore the first few softmax predictions in the distillation loss for each sample. We evaluated the utility of each method by measuring the perplexity (PPL) of the test set.
@@ 5.1 Experimental Setup
5.2 Main Results
Table 2 shows the results. We see that for the zero-shot performance of the student model, i.e. $\epsilon=0$, which is just a pre-trained DistilGPT2 model, it achieves a perplexity score of 65.46 and 57.42 on Yelp and Big Patent, respectively. After performing DP-SGD with $\epsilon=2$, we see that the student improves to 48.12 and 41.80, which is a substantial utility improvement over the pre-trained student. The DP-SGD baseline, directly privately fine-tuning the student, serves as a lower bound for the performance of the student. Meaning that our framework must perform at least better than DP-SGD, or else the knowledge distilled from the teacher only hurts the performance of the student.
For the DPKD baseline, we evenly allocated the privacy budget $\epsilon$ between the teacher and student, i.e., $T$ and $S$ use $\epsilon=1$ for DP-SGD. We observe that DPKD improves over the DP-SGD baseline by 2 and 0.6 PPL on the Yelp and Big Patent datasets, respectively. Hence, DPKD can give a significant performance boost to the student models compared to directly privately fine-tuning them for generative tasks, which was not the case for classification tasks shown in Mireshghallah et al. (2022).
Our method, using a GPT2-Large model with $\epsilon=2$ as the teacher, was able to substantially outperform all baselines for both datasets. For the Yelp datasets, our method beats all baselines by at least 2 PPL. The margin is much larger for the Big Patent dataset, as our method outperforms all baselines by at least nearly 9 PPL. One contributing factor for the margin being larger for Big Patent is that the number of DP synthetic data samples that were generated is larger than the size of the original dataset by an additional 50K data samples. Most importantly, the results for both our method and DPKD demonstrate that the traditional knowledge distillation of aligning outputs between the teacher and the student can boost the performance of the student for differentially private generative tasks.
One insightful observation is that DP Syn Data severely lags behind all baselines, only providing a modest improvement over the zero-shot baseline but not improving over the DP-SGD baseline. Additionally, we found that it is harder to optimize DP Syn Data due to the student easily overfitting the synthetic review data. This highlights the fact that training on DP synthetically generated data alone is insufficient to competitively train a compressed DP student model for generative tasks. We surmise that a teacher model at the scale of a large foundation model is needed to generate high enough quality synthetic text to overcome this limitation Yu et al. (2023a).
@@ 5.3 Ablation Study
Figure 1(a) plots the change of the PPL score as $\lambda$ changes. We see that relying purely on the distillation loss $\lambda=1$ already achieves competitive performance, capturing most of the knowledge from the private dataset. Meanwhile, the best utility we could achieve when fine-tuning only on the synthetic data, $\lambda=0$, is a severe utility degradation compared to $\lambda=1$. Interestingly, nearly evenly combining both losses, $\lambda=0.4$, leads to the most optimal performance. Hence both the supervised learning and distillation loss in Equation 1 are needed.
@@ 5.4 Exploiting Hidden Representation of The Teacher
The main results in section 5.2 showed that there is still a reasonable utility gap between the PPL score of the DP teacher model with $\epsilon=1$ and our student model with $\epsilon=2$. The difference for the Yelp dataset is 17.3 perplexity, while the difference for Big Patent is less substantial, about 10.5 perplexity. In this section, we attempt to add more ways for the teacher to distill knowledge onto the student via a Mean Squared Error (MSE) loss $\alpha\mathcal{L}_{\text{MSE}}$ on the hidden representations to Eq 1. The rationale behind this method is that it would be beneficial to align the hidden representations of the student with the hidden representations of the teacher, assuming that the teacher has better hidden representations. Hence, we can use the MSE loss to push the student to mimic the hidden representation of the teacher. Note that since $T$ was fine-tuned using DP-SGD, by the post-processing property of DP, adding the MSE loss to Eq 1 will not leak additional privacy.
Table 3 compares the performance between the teacher and the student. We see that the utility difference between the teacher and the student is even smaller, only 6 PPL. Table 4 shows that the inclusion of the MSE loss into Eq 1 does mildly improve the overall utility of the student. We suspect that the distillation loss $\mathcal{L}_{\text{KL}}$ already captures most of the information distilled from the teacher. However, the improvement from the MSE loss suggests that our framework can be further improved by including more information about the hidden representation of the teacher. We leave it as a future work to explore other advanced knowledge distillation techniques that exploit this information.
@@
Table 3: Further reducing the utility gap between the teacher and the student by adding an MSE loss to Eq 1. The results are for $\epsilon=2$ and use the Big Patent dataset.
@@ 6 Conclusion
In this work, we introduced a novel DP knowledge distillation algorithm by utilizing DP synthetic data generation. We presented comprehensive results that show our framework obtains the best utility over all existing baselines for strict privacy settings, and we proved that differentially private knowledge distillation for generative large language models can be effective. The ablation study provided shows that both the synthetic text as the hard label and the output distribution of the teacher as the soft label are crucial to boosting the performance of the student. We can further improve the main results of our framework by generative more synthetic text data and including more knowledge distillation of the hidden representations.
@@ 7 Limitations
One limitation of our work is that the teacher model needs to be trained with DP-SGD. This is because the teacher is a function of the private dataset, and is used for synthesizing text and aligning its outputs with the output of the student. However, the teacher is generally much larger than the student, and thus the computational and memory costs of running DP-SGD for the teacher are much larger than for the student. Furthermore, the problem setup of DP compression only requires differential privacy for the student, not for the teacher. A more optimal solution to this problem would only use DP-SGD for the student, and avoid applying DP-SGD for the teacher. However, such a solution is non-trivial to derive and is left as future work.
@@ 8 Ethical Considerations
In this work, we utilized pre-trained large language models and well-known language modeling datasets that were accessed from the Hugging Face API, which are publicly available and free to use. The Big Patent dataset is licensed under CC BY 4.0. The GPT2- models are licensed under the Apache License, Version 2.0. We could not find a license for the Yelp dataset, however, the intended use of this dataset must be for academic research, which we do follow. Our intended use of the artifacts in this work is academic in nature, which is aligned with the intended use of the creators. We did not anonymize the dataset since Big Patent does not contain personally identifiable information, and we removed attributes that contain any private information, such as user or business id, from the Yelp dataset, as the dataset we worked with only contained text data and non-private categorical data, which are used as control codes. Additionally, we did not disclose the contents of either dataset. Our use of public models and datasets minimizes any unintended privacy leakage that could result from experimenting.

### v2

@@ 4.3 Privacy Analysis
Proof. Our goal is for DistilDP to guarantee $(\epsilon,\delta)$-DP for $D$, which composes of text content for each data sample. After step 1, our method produces a teacher model $T$ that is $(\epsilon,\delta)$-DP. In step 2, sampling from $T$ incurs no extra privacy loss due to the post-processing property. However, one subtle, but crucial, detail is the potential privacy leakage from the control codes $C$. We limit the control codes to only contain categorical information of a data sample, not text contents. However, the categorical distribution in the original dataset may be private information. In our work, we follow the same procedure as Yue et al. (2022) by ignoring this low-cost privacy loss and using the exact categorical distribution. For step 3, since $\tilde{D}$ and $T$ are $(\epsilon,\delta)$-DP with respect to $D$, utilizing a non-private optimizer to minimize Equation 3.1 will result in $S$ being $(\epsilon,\delta)$-DP due to the post-processing property. ∎
Parameter DP-SGD DPKD DP Syn Ours
Teacher
Epochs - 101, 202, 3 101, 202, 3 101, 202, 3
Learning Rate - 1e-41, 2, 3 1e-41, 2, 3 1e-41, 2, 3
Batch Size - 40961, 2, 3 40961, 2, 3 40961, 2, 3
Clipping Norm - 1.01, 2, 3 1.01, 2, 3 1.01, 2
Student
Epochs 251, 402, 3 251, 402, 3 51, 32, 3 121, 2, 3
Learning Rate 1e-41, 2, 3 1e-41, 2, 3 1e-61, 2e-52, 3 8e-51, 2, 3
Batch Size 40961, 2, 3 40961, 2, 3 161, 2, 3 161, 2, 3
Clipping Norm 1.01, 2, 3 1.01, 2, 3 - -
Table 1: Training hyperparameters for the teacher and student model of every method for the 1Yelp, 2Big Patent, and 3DBpedia datasets.
@@ 5 Experiments
5.1 Experimental Setup
We experimentally evaluated the privacy-utility tradeoff of our framework with the following setup. We used GPT2-Large Radford et al. (2019) and DistilGPT2 Sanh et al. (2019) as our teacher and student model, respectively. GPT2-Large contains 774M parameters while DistilGPT2 contains 82M parameters, so the teacher model is about 9.5 times larger than the student. Our implementation of fine-tuning language models with DP and knowledge distillation utilizes the dp-transformers library Inan et al. (2022), which combines the NLP library Huggingface Wolf et al. (2019) with the DP library Opacus Yousefpour et al. (2021).
We experimented with the Yelp Open dataset  [^1] https://www.yelp.com/dataset, which contains review text data on businesses. We follow the experimental setup as Yue et al. (2022) by using the review stars and business category as the control codes. These control codes are constructed as "Business Type: Restaurant |||| Review Stars: 3.0" and are prepended to each sample. The top 10 frequent business categories are sampled and reviews with no ratings are removed. The resulting data split is $1.9$M reviews for training, $5000$ for validation, and $5000$ for testing. Each data sample is set to a sequence length of $128$.
Lastly, we experimented on the DBpedia dataset Zhang et al. (2015), which is an ontology classification dataset consisting of 14 non-overlapping classes from DBpedia 2014. Each of these 14 ontology classes are used as a control code for the training samples. In total, we randomly selected 195K training samples, 5000 samples for validation, and 5000 samples for the test set.
We compare our method against three baselines: (1) a student model fine-tuned directly with DP-SGD, namely the teacher plays no role in training the student; (2) a student model fine-tuned using DPKD Mireshghallah et al. (2022); (3) a student model fine-tuned only on DP synthetic data (DP Syn Data), i.e. $\lambda=0$ in Eq. 3.1, which follows the framework of Yu et al. (2023a). The training hyperparameters used for our method and the baselines with the teacher and student models are shown in table 1. We set the privacy budget to $\epsilon=2$, which is a strong privacy guarantee. We follow the standard practice of setting $\delta=1/N$ where $N$ is the size of the dataset.
@@ 5.1 Experimental Setup
GPT2-Large DP-SGD 1.0 - 26.84 21.91 28.85
DistilGPT2 Zero-shot 0 - 65.46 57.42 83.00
DistilGPT2 DP-SGD 2.0 - 48.12 41.80 60.81
DistilGPT2 DPKD 2.0 GPT2-Large 46.16 41.20 59.85
DistilGPT2 DP Syn Data 2.0 GPT2-Large 61.84 49.26 59.87
DistilGPT2 DistilDP 2.0 GPT2-Large 44.15 32.43 49.11
@@
For the generation of synthetic text, we use the truncation decoding strategies of top-$k$ Fan et al. (2018) and top-$p$ Holtzman et al. (2019) sampling using $k=50$ and $p=0.9$. The teacher model generates 400K synthetic samples using control codes from the Yelp and DBpedia training dataset and 200K synthetic samples from Big Patent. The maximum sequence length of each generated sequence is $128$. We follow the methodology of Kurakin et al. (2023) for training on DP synthetic data by splitting the Yelp and DBPedia synthetic samples into $396$K for the training set and $4$K for the validation. Likewise, $198$K and $2$K Big Patent synthetic examples for the training and validation set, respectively. For both our method and DPKD, we set $t=1.0$ and $\lambda=0.4$ for Equation 3.1. One minor technicality of our method is that because our teacher model was trained on control codes, its output distribution on the initial context vectors will be skewed toward the types of control codes. Hence, we ignore the first few softmax predictions in the distillation loss for each sample. We evaluated the utility of each method by measuring the perplexity (PPL) of the test set.
5.2 Main Results
Table 2 shows the results. We see that for the zero-shot performance of the student model, i.e. $\epsilon=0$, which is just a pre-trained DistilGPT2 model, it achieves a perplexity score of $65.46$, $57.42$, and $83.00$ on Yelp, Big Patent, and DBpedia, respectively. After performing DP-SGD with $\epsilon=2$, we observe that the student utility improves to $48.12$, $41.80$, and $60.81$, a significant utility improvement over the pre-trained student. However, compared to the DP fine-tuned GPT2-Large model with $\epsilon=1$, there is still much improvement to strive for.
The DP-SGD baseline, where the student is privately fine-tuned without any teacher, serves as a lower bound for the utility of the student. This means that our framework must perform at least better than DP-SGD, or else the knowledge distilled from the teacher only hurts the student’s utility.
For the DPKD baseline, we evenly allocated the privacy budget $\epsilon$ between the teacher and student, i.e., $T$ and $S$ use $\epsilon=1$ for DP-SGD. We observe that DPKD improves over the DP-SGD baseline by $2$, $0.6$ and $0.96$ PPL on the Yelp, Big Patent, DBpedia datasets, respectively. Hence, DPKD can give a modest performance boost to the student models compared to directly privately fine-tuning them for generative tasks, which was not the case for classification tasks shown in Mireshghallah et al. (2022).
@@ 5.2 Main Results
DistilDP, using a GPT2-Large model with $\epsilon=2$ as the teacher, was able to substantially outperform all baselines for both datasets. For the Yelp datasets, our method beats all baselines by at least $2$ PPL. The margin is much larger for the Big Patent and DBpedia dataset, as our method outperforms all baselines by at least $9$ PPL. One contributing factor for the margin being larger for Big Patent and DBPedia is that the difference in size between the synthetic and original dataset is smaller than for Yelp. Most importantly, the results for both our method and DPKD demonstrate that the traditional knowledge distillation of aligning outputs between the teacher and the student can boost the performance of the student for differentially private generative tasks.
One insightful observation is that for the Yelp and Big Patent datasets, DP Syn Data severely lags behind all baselines, only providing a modest improvement over the zero-shot baseline but not improving over the DP-SGD baseline. However, we do observe that for DBpedia, DP Syn Data can match the utility of DPKD. Additionally, we found that it is harder to optimize DP Syn Data due to the student easily overfitting the synthetic review data. This result highlights the fact that training on DP synthetically generated data alone is insufficient to competitively train a compressed DP student model for generative tasks. We surmise that a teacher model at the scale of a large foundation model is needed to generate high enough quality synthetic text to overcome this limitation Yu et al. (2023a).
@@ 5.3 Ablation Study
Figure 2(a) plots the change of the PPL score as $\lambda$ changes. We see that relying purely on the distillation loss $\lambda=1$ already achieves competitive performance, capturing most of the knowledge from the private dataset. Meanwhile, the best utility we could achieve when fine-tuning only on the synthetic data, $\lambda=0$, is a severe utility degradation compared to $\lambda=1$. Interestingly, nearly evenly combining both losses, $\lambda=0.4$, leads to the most optimal performance, necessitating both the supervised learning and distillation loss in Equation 3.1.
@@ 5.4 Exploiting Hidden Representation of The Teacher
The main results in section 5.2 showed that there is still a reasonable utility gap between the PPL score of the DP teacher model with $\epsilon=1$ and the DistilDP student model with $\epsilon=2$. The difference for the Yelp dataset is 17.3 perplexity, while the difference for Big Patent is less substantial, about 10.5 perplexity. In this section, we attempt to add more ways for the teacher to distill knowledge onto the student via a Mean Squared Error (MSE) loss $\alpha\mathcal{L}_{\text{MSE}}$ on the hidden representations to Eq 3.1. The rationale behind this method is that it would be beneficial to align the hidden representations of the student with the hidden representations of the teacher, assuming that the teacher has better hidden representations. Hence, we can use the MSE loss to push the student to mimic the hidden representation of the teacher. Note that since $T$ was fine-tuned using DP-SGD, by the post-processing property of DP, adding the MSE loss to Eq 3.1 will not leak additional privacy.
@@
Table 3: Further reducing the utility gap between the teacher and the student by adding an MSE loss to Eq 3.1. The results are for $\epsilon=2$ and use the Big Patent dataset.
@@
Table 3 compares the performance between the teacher and the student. We see that the utility difference between the teacher and the student is even smaller, only 6 PPL. Table 4 shows that the inclusion of the MSE loss into Eq 3.1 does mildly improve the overall utility of the student. We suspect that the distillation loss $\mathcal{L}_{\text{KL}}$ already captures most of the information distilled from the teacher. However, the improvement from the MSE loss suggests that our framework can be further improved by including more information about the hidden representation of the teacher. We leave it as a future work to explore other advanced knowledge distillation techniques that exploit this information.
@@ 6 Conclusion
In this work, we introduced DistilDP, a novel DP knowledge distillation algorithm by utilizing DP synthetic data generation. We presented comprehensive experimental results that demonstrate DistilDP obtains the best utility over all existing baselines for strict privacy settings. More generally, we showed that differentially private knowledge distillation for autoregressive large language models can be effective. The provided ablation study highlights that both the synthetic text as the hard label and the output distribution of the teacher as the soft label are crucial to boosting the performance of the student. We believe further improvements to our framework are feasible by generating more synthetic text data and including additional knowledge distillation from the hidden representations.
@@ 7 Limitations
One limitation of our work is that the teacher model needs to be trained with DP-SGD. This limitation is because the teacher is a function of the private dataset, and is used for synthesizing text and aligning its outputs with the output of the student. However, the teacher is generally much larger than the student, and thus the computational and memory costs of running DP-SGD for the teacher are much larger than for the student. Furthermore, the problem setup of DP compression only requires differential privacy for the student, not for the teacher. A more optimal solution to this problem would only use DP-SGD for the student, and avoid applying DP-SGD for the teacher. However, such a solution is non-trivial to derive and is left as future work.
@@ 8 Ethical Considerations
In this work, we utilized pre-trained LLMs and well-known language modeling datasets that were accessed from the Hugging Face API, which are publicly available and free to use. The Big Patent dataset is licensed under CC BY 4.0. The GPT2- models are licensed under the Apache License, Version 2.0. We could not find a license for the Yelp dataset, however, the intended use of this dataset must be for academic research, which we do follow. Our intended use of the artifacts in this work is academic in nature, which is aligned with the intended use of the creators. We did not anonymize the dataset since Big Patent does not contain personally identifiable information, and we removed attributes that contain any private information, such as user or business id, from the Yelp dataset, as the dataset we worked with only contained text data and non-private categorical data, which are used as control codes. Additionally, we did not disclose the contents of either dataset. Our use of public models and datasets minimizes any unintended privacy leakage that could result from experimenting.
Acknowledgements
This work is supported by the Defense Advanced Research Projects Agency (DARPA) under Contract Nos. HR001120C0088, NSF award number 2224319 and DGE-1842487, REAL@USC-Meta center, and gifts from VMware. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.